### v1

# A HHO formulation for variable density incompressible flows where the density is purely advected

Abstract
We propose a Hybrid High-Order (HHO) formulation of the incompressible Navier–Stokes equations with variable density that provides exact conservation of volume and, accordingly, pure advection of the density variable. The spatial discretization relies on hybrid velocity-density-pressure spaces and the temporal discretization is based on Explicit Singly Diagonal Implicit Runge-Kutta (ESDIRK) methods. The formulation possesses some attractive features that can be fruitfully exploited for the simulation of mixtures of immiscible incompressible fluids, namely: pressure-robustness, conservation of volume enforced cell-by-cell up to machine precision, robustness in the convection dominated regime thanks to upwind stabilizations, weak imposition of boundary conditions, implicit high-order accurate time stepping encompassing unsteady boundary conditions, reduced memory footprint thanks to static condensation, possibility to exploit inherited $p$-multilevel solution strategies to improve performance of iterative solvers. Numerical validation is performed showcasing convergence rates, pressure-robustness and pure advection of the density, thus we tackle the Rayleigh-Taylor instability at Reynolds 10001~000 and 50005~000.
1 Introduction
In recent years, the introduction of Hybrid High-Order (HHO) and Hybridizable Discontinuous Galerkin (HDG) methods has provided new lifeblood to promote the development of high-order accurate computational modelling tools for incompressible fluid flows. In contrast to Discontinuous Galerkin (DG) methods, HHO and HDG methods are based on degrees of freedom that are broken polynomials on both the mesh and its skeleton (that is, the set collecting interelement boundaries). Relevant features of hybrid schemes are: i) local (cell-by-cell) conservation of physical quantities, ii) increased convergence rates in the diffusion-dominated regime iii) availability of upwind stabilizations to improve the robustness in the convection-dominated regime iv) resiliency to mesh distortion and grading, v) implicit time integration with reduced memory footprint of the Jacobian matrix.
In this work we extend the HHO formulations of [7]-[8], devoted to constant density incompressible flows, to cope with mixtures of immiscible incompressible fluids with variable density. The present research effort relies on a hybrid discretization of all the variables of the problem, including velocity, pressure and density. The choice of a hybrid pressure, combined with an appropriate selection of local polynomial spaces, allows to reconstruct an $\bm{H}(\operatorname{div})$-conforming velocity approximation starting from non-conforming polynomial spaces. Thanks to $\bm{H}(\operatorname{div})$-conformity and exact incompressibility, the mass conservation equation can be formulated as the material time derivative of the density, both at the continuous and at the discrete level. Moreover, the HHO discretization of mass transport uses to good advantage the availability of a solenoidal velocity field, see e.g., Di Pietro et al. [13]. Pressure-robustness implies that irrotational body forces only affect the pressure field and, as a consequence, the velocity and density errors are insensitive to the accuracy of the pressure field representation. Robustness in the convection-dominated regime is achieved by introducing upwinding stabilizations for the transport of mass and momentum, in tandem with suitable discretization of the density time derivative.
In order to treat time-dependent problems, we rely on high-order accurate Explicit Singly Diagonal Implicit Runge-Kutta (ESDIRK) time integration schemes especially conceived for Differential Algebraic Equations (DAEs) of differential index equal to two. Efficient resolution of the algebraic problems resulting from the space-time discretization relies on $p$-multigrid preconditioners in the spirit of Botti et al. [9], specifically conceived to work in tandem with Schur complement solution strategies. Thanks to static condensations, the blocked structure, the dimension of the blocks and, ultimately, the sparsity of the global matrix resulting from HHO discretizations is dictated by face unknowns. Interestingly, since the dimension of d1d{-}1-variate polynomial spaces grows slower that the dimension of $d$-variate polynomial spaces when increasing $k$, both the memory footprint of the global matrix and the cost of matrix-vector products involved in iterative solution strategies benefit from static condensation.
Several methods for variable density incompressible flows have been proposes in literature. A DG formulation based on exact Riemann solvers for the artificial compressibility perturbation of the problem was proposed by Bassi et al. [3] and assessed in [2]. An entropy–stable DG approximation was devised by Manzanero et al. [21] while a post-processed semi-implicit DG scheme was introduced by [20]. A Discrete Duality Finite Volume (DDFV) scheme was proposed by Goudon et al. [15]. Recently, Fu [14] addressed the Cahn-Hilliard phase-field model with a HDG formulation based on divergence-free velocity spaces. Earlier works focused on projection methods, consider for example Guermond et al. [16]-[17].
The paper is organized as follows. In Sections 2 and 3 we respectively introduce the model problem and present its discrete formulation. $\bm{H}(\operatorname{div})$-conformity and pressure robustness are addressed in Section 3.5. Numerical validation of spatial and temporal convergence rates is performed in Section 4.1 and Section 4.2 considering exact and manufactured solutions. A realistic flow configuration is tackled in Section 4.3 where we consider a low Atwood number Rayleigh-Taylor instability at $\mathrm{Re}{=}1~000$ and $\mathrm{Re}{=}5~000$. Finally, some conclusions are drawn in Section 5.
2 Continuous setting
Given a polygonal or polyhedral domain $\Omega\subset\mathbb{R}^{d}$, $d\in\{2,3\}$, with boundary $\partial\Omega$, the initial divergence-free velocity field $\bm{u}_{0}:\Omega\rightarrow\mathbb{R}^{d}$, the initial density field $\rho_{0}:\Omega\rightarrow\mathbb{R}^{+}$ and a finite time $t_{\rm F}$, the variable density incompressible Navier–Stokes problem consists in finding the velocity field $\bm{u}:\Omega\times[0,t_{\rm F}]\rightarrow\mathbb{R}^{d}$, the density field $\rho:\Omega\times(0,t_{\rm F}]\rightarrow\mathbb{R}^{+}$, and the pressure field $p:\Omega\times(0,t_{\rm F}]\rightarrow\mathbb{R}$, such that $\bm{u}(\cdot,0)=\bm{u}_{0}$, $\rho(\cdot,0)=\rho_{0}$ and
$\frac{\partial(\rho\bm{u})}{\partial t}+\nabla\cdot\left[{(\rho\bm{u}\otimes\bm{u})-\mu\nabla\bm{u}}\right]+\nabla p$ $=\bm{f}$ $\text{in $\Omega\times(0,t_{\rm F}]$},$ (1a)
$\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\bm{u})$ $=0$ $\text{in $\Omega\times(0,t_{\rm F}]$},$ (1b)
$\nabla\cdot\bm{u}$ $=0$ $\text{in $\Omega\times(0,t_{\rm F}]$},$ (1c)
$\bm{u}$ $=\bm{g}_{\rm D}$ $\text{on $\partial\Omega_{\rm D}\times(0,t_{\rm F}]$},$ (1d)
$p\bm{n}-\mu(\nabla\bm{u})\bm{n}$ $=\bm{g}_{\rm N}$ $\text{on $\partial\Omega_{\rm N}\times(0,t_{\rm F}]$},$ (1e)
[…]
where $\bm{n}$ denotes the unit vector normal to $\partial\Omega$ pointing out of $\Omega$, $\mu>0$ is the (constant) viscosity, $\bm{g}_{\rm D}$ and $\bm{g}_{\rm N}$ denote, respectively, the prescribed velocity on the Dirichlet boundary $\partial\Omega_{\rm D}\subset\partial\Omega$ and the prescribed traction on the Neumann boundary $\partial\Omega_{\rm N}\coloneqq\partial\Omega\setminus\partial\Omega_{\rm D}$, while $\bm{f}:\Omega\to\mathbb{R}^{d}$ is a given body force. Since, by virtue of the incompressibility constraint, we consider the Laplace operator in place of the divergence of the symmetric gradient, tractions do not account for tangential variations of the normal velocity component. The density is prescribed on inflow boundaries, defined as follows
$\partial\Omega_{{\rm In}}\coloneqq\left\{\bm{x}\in\partial\Omega\;\text{such that}\;\left\{\begin{array}[]{lcl}\bm{g}_{\rm D}\cdot\bm{n}<0&\text{if}&\bm{x}\in\partial\Omega_{\rm D},\\ \bm{u}\cdot\bm{n}<0&\text{if}&\bm{x}\in\partial\Omega_{\rm N}\end{array}\right.\right\}$.
Equations (1a), (1b) and (1c) are the linear momentum, the mass conservation and the volume conservation equations, respectively. We remark that the discrete formulation will be designed mimicking as closely as possible the weak formulation of the continuous problem. Accordingly, conservation of volume will be exactly enforced at the discrete level, ensuring that the density is a purely advected quantity.
3 ESDIRK-HHO discretization
3.1 Mesh
The HHO formulation proposed in this work is capable of dealing with two and three dimensional flow problems. Nevertheless, since numerical test cases focus on two space dimensions, discrete settings are provided for the 2D case focusing on simplicial meshes. We consider meshes of the domain $\Omega$ corresponding to couples $\mathcal{M}_{h}\coloneq(\mathcal{T}_{h},\mathcal{F}_{h})$, where $\mathcal{T}_{h}$ is a finite collection of triangular elements such that $h\coloneq\max_{T\in\mathcal{T}_{h}}h_{T}>0$ with $h_{T}$ denoting the diameter of $T$, while $\mathcal{F}_{h}$ is a finite collection of line segments, with $h_{F}$ denoting the length of $F$. It is assumed henceforth that the mesh $\mathcal{M}_{h}$ is shape- and contact-regular, as detailed in [12, Definition 1.4] and that its trace on $\partial\Omega$ is compatible with the partition $\partial\Omega=\partial\Omega_{\rm D}\sqcup\partial\Omega_{\rm N}$. We respectively denote by $\mathcal{F}_{h}^{\rm D}$ and $\mathcal{F}_{h}^{\rm N}$ the sets of Dirichlet and Neumann faces, whereas the set of internal faces is $\mathcal{F}_{h}^{\rm i}\coloneq\mathcal{F}_{h}\setminus\big(\mathcal{F}_{h}^{\rm D}\cup\mathcal{F}_{h}^{\rm N}\big)$. For each mesh cell $T\in\mathcal{T}_{h}$, the faces contained in the cell boundary $\partial T$ are collected in the set $\mathcal{F}_{T}$ and we additionally let $\partial T^{\bullet}\coloneqq\partial T\cap\partial\Omega_{\bullet}$ for $\bullet\in\{{\rm D},{\rm N}\}$. For all $T\in\mathcal{T}_{h}$ and all $F\in\mathcal{F}_{T}$, $\bm{n}_{TF}$ denotes the unit vector normal to $F$ pointing out of $T$ and $\bm{n}_{\partial T}:\partial T\to\mathbb{R}^{d}$ is such that $\bm{n}_{\partial T}|_{F}\coloneqq\bm{n}_{TF}$ for all $F\in\mathcal{F}_{T}$.
3.2 Discrete spaces
Hybrid High-Order methods hinge on local polynomial spaces on mesh cells and faces. For given integers $\ell\geq 0$ and $n\geq 1$, we denote by $\mathbb{P}_{n}^{\ell}$ the space of $n$-variate polynomials of total degree $\leq\ell$ (in short, of degree $\ell$). For $X$ mesh cell or face, we denote by $\mathcal{P}^{\ell}(X)$ the space spanned by the restriction to $X$ of functions in $\mathbb{P}_{d}^{\ell}$.
The global discrete spaces for the velocity, density and pressure unknowns are respectively defined as follows:
[…]
$\underline{Z}_{h}^{k}\coloneqq\left\{\underline{z}_{h}=\big((z_{T})_{T\in\mathcal{T}_{h}},(z_{F})_{F\in\mathcal{F}_{h}}\big)\,:\,\begin{array}[]{l}\text{$z_{T}\in\mathcal{P}^{k}(T)$ for all $T\in\mathcal{T}_{h}$},\\ \text{$z_{F}\in\mathcal{P}^{k}(F)$ for all $F\in\mathcal{F}_{h}$}\end{array}\right\},$
$\underline{Q}_{h}^{k+1}\coloneqq\left\{\underline{q}_{h}=\big((q_{T})_{T\in\mathcal{T}_{h}},(q_{F})_{F\in\mathcal{F}_{h}}\big)\,:\,\begin{array}[]{l}\text{$q_{T}\in\mathcal{P}^{k}(T)$ for all $T\in\mathcal{T}_{h}$},\\ \text{$q_{F}\in\mathcal{P}^{k+1}(F)$ for all $F\in\mathcal{F}_{h}$}\end{array}\right\}.$
The restriction of these spaces and their elements to a mesh cell $T\in\mathcal{T}_{h}$ are denoted replacing the subscript $h$ by $T$ and are obtained collecting the components attached to $T$ and its boundary. Given $\underline{\bm{v}}_{T}\in\underline{\bm{V}}_{T}^{k}$, we will also denote by $\bm{v}_{\partial T}$ the function such that $\bm{v}_{\partial T}|_{F}=\bm{v}_{F}$ for all $F\in\mathcal{F}_{T}$. Given $\underline{q}_{T}\in\underline{Q}_{T}^{k+1}$ and $\underline{z}_{T}\in\underline{Z}_{T}^{k}$, the notation $q_{\partial T}$ and $z_{\partial T}$ are defined similarly.
3.3 Gradient reconstruction operators
For each mesh cell $T\in\mathcal{T}_{h}$, the tensor and vector gradient reconstruction operators ${\mathcal{G}}_{T}^{k}:\underline{\bm{V}}_{T}^{k}\to\mathcal{P}^{k}(T)^{d\times d}$ and ${\bm{\mathfrak{g}}}_{T}^{k+1}:\underline{Q}_{T}^{k+1}\to\mathcal{P}^{k+1}(T)^{d}$ are respectively defined as follows: For all $(\underline{\bm{v}}_{T},\underline{q}_{T})\in\underline{\bm{V}}_{T}^{k}\times\underline{Q}_{T}^{k+1}$,
$\int_{T}{\mathcal{G}}_{T}^{k}\underline{\bm{v}}_{T}:\bm{\tau}=\int_{T}\nabla\bm{v}_{T}:\bm{\tau}-\int_{\partial T}(\bm{v}_{T}-\bm{v}_{\partial T})\otimes\bm{n}_{\partial T}:\bm{\tau}\qquad\forall\bm{\tau}\in\mathcal{P}^{k}(T)^{d\times d},$
$\int_{T}{\bm{\mathfrak{g}}}_{T}^{k+1}{\underline{q}_{T}}\cdot{\bm{v}}=-\int_{T}{q}_{T}\;\nabla\cdot{\bm{v}}+\int_{\partial T}{q}_{\partial T}{\bm{v}}\cdot\bm{n}_{\partial T}\qquad\forall{\bm{v}}\in\mathcal{P}^{k+1}(T)^{d}.$
3.4 Semi-discrete HHO formulation
We first present the semi-discrete formulation of the variable density problem in (1) omitting the terms that are responsible of the weak imposition of boundary conditions. Those will be introduced thereafter for the sake of readability.
Given $(\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T})\in\underline{\bm{V}}_{T}^{k}\times\underline{Z}_{T}^{k}\times\underline{Q}_{T}^{k+1}$, the (local) spatial discretizations $sd^{\rm slm}_{T,{\rm!bc}}((\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T});\cdot):\underline{\bm{V}}_{T}^{k}\to\mathbb{R}$ of the steady momentum equation, $sd^{\rm sms}_{T,{\rm!bc}}(\underline{\bm{u}}_{T},\underline{\rho}_{T};\cdot):\underline{Z}^{k}_{T}\to\mathbb{R}$ of the steady mass conservation equation and $sd^{\rm vol}_{T,{\rm!bc}}(\underline{\bm{u}}_{T};\cdot):\underline{Q}^{k+1}_{T}\to\mathbb{R}$ of the volume conservation equation are such that, for all $\underline{\bm{v}}_{T}\in\underline{\bm{V}}_{T}^{k}$, all $\underline{z}_{T}\in\underline{Z}_{T}^{k}$ and all $\underline{q}_{T}\in\underline{Q}_{T}^{k+1}$,
$sd^{\rm slm}_{T,{\rm!bc}}$ $\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T});\underline{\bm{v}}_{T}\bigr)=-\int_{T}\rho_{T}(\bm{u}_{T}\otimes\bm{u}_{T}):\nabla\bm{\pi}_{T}^{k}\bm{v}_{T}$ (2)
$+\int_{\partial T}\left[\bigl(\bm{u}_{T}{\cdot}\bm{n}_{\partial T}\bigr)^{\oplus}{\rho}_{T}\bm{u}_{T}{+}\bigl(\bm{u}_{T}{\cdot}\bm{n}_{\partial T}\bigr)^{\ominus}\rho_{\partial T}\bm{u}_{\partial T}\right]\cdot(\bm{\pi}_{T}^{k}\bm{v}_{T}-\bm{\pi}_{\partial T}^{k}\bm{v}_{\partial T})$ (3)
$+\int_{T}\mu\,{{\mathcal{G}}_{T}^{k}}\underline{\bm{u}}_{T}:{{\mathcal{G}}_{T}^{k}}\underline{\bm{v}}_{T}+\int_{\partial T}\frac{\mu}{h_{T}}\,\bm{\pi}_{\partial T}^{k}(\bm{u}_{T}-\bm{u}_{\partial T})\cdot\bm{\pi}_{\partial T}^{k}(\bm{v}_{T}-\bm{v}_{\partial T})$ (4)
$+\int_{T}{\bm{\mathfrak{g}}}_{T}^{k+1}\underline{p}_{T}\cdot\bm{v}_{T}-\int_{T}\bm{f}\cdot\bm{v}_{T},$ (5)
$sd^{\rm sms}_{T,{\rm!bc}}$ $\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T});\underline{z}_{T}\bigr)=$
$-\int_{T}\rho_{T}\bm{u}_{T}\cdot\nabla z_{T}\,+\int_{\partial T}\left[\bigl(\bm{u}_{T}{\cdot}\bm{n}_{TF}\bigr)^{\oplus}{\rho}_{T}{+}\bigl(\bm{u}_{T}{\cdot}\bm{n}_{TF}\bigr)^{\ominus}\rho_{\partial T}\right]({z}_{T}-{z}_{\partial T}),$ (6)
$sd^{\rm vol}_{T,{\rm!bc}}$ $(\underline{\bm{u}}_{T};\underline{q}_{T})=\int_{T}{\bm{\mathfrak{g}}}_{T}^{k+1}\underline{q}_{T}\cdot\bm{u}_{T},$ (7)
where $\alpha^{\oplus}\coloneqq\frac{1}{2}(\alpha+|\alpha|)$ and $\alpha^{\ominus}\coloneqq\frac{1}{2}(\alpha-|\alpha|)$, $\bm{\pi}_{\partial T}^{k}$ is the $L^{2}$-orthogonal projector onto the broken polynomial space $\mathcal{P}^{k}(\partial T)^{d}$ defined as follows
$\mathcal{P}^{k}(\partial T)^{d}\coloneqq\left\{\bm{v}\in L^{2}(\partial T)^{d}\;:\;\text{$\bm{v}|_{F}\in\mathcal{P}^{k}(F)^{d}$ for all $F\in\mathcal{F}_{T}$}\right\},$
and $\bm{\pi}_{T}^{k}$ is the $L^{2}$-orthogonal projector onto $\mathcal{P}^{k}(T)^{d}$. We remark that upwinding is introduced in the mass conservation and linear momentum equations for enhancing the stability of the discrete formulation. The idea of utilizing $L^{2}$-projections of test functions in the convective term formulation of (2)-(3) is inspired by [4] and improves the accuracy of the method in the convection-dominated regime. The diffusion term stabilization of (4), also involving $L^{2}$-projection operators, was first proposed by [19], see also [11]. Note that the HHO discretization of the mass conservation equation in (6) only relies on the cell velocity $\bm{u}_{T}$ which, as result of the HHO discretization of the volume conservation in (7), is $\bm{H}(\operatorname{div})$-conforming and divergence-free, see Section 3.5 for details. Accordingly, the trace of the normal velocity component is continuous on $\partial T$.
Let’s now take into account Dirichlet and Neumann boundary conditions. Given $(\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T})\in\underline{\bm{V}}_{T}^{k}\times\underline{Z}_{T}^{k}\times\underline{Q}_{T}^{k+1}$, the (local) spatial discretizations $sd^{\rm slm}_{T}((\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T});\cdot):\underline{\bm{V}}_{T}^{k}\to\mathbb{R}$ of the steady momentum equation, $sd^{\rm sms}_{T}(\underline{\bm{u}}_{T},\underline{\rho}_{T};\cdot):\underline{Z}^{k}_{T}\to\mathbb{R}$ of the steady mass conservation equation and $sd^{\rm vol}_{T}(\underline{\bm{u}}_{T};\cdot):\underline{Q}^{k+1}_{T}\to\mathbb{R}$ of the volume conservation equation are such that, for all $\underline{\bm{v}}_{T}\in\underline{\bm{V}}_{T}^{k}$, all $\underline{z}_{T}\in\underline{Z}_{T}^{k}$ and all $\underline{q}_{T}\in\underline{Q}_{T}^{k+1}$,
$sd^{{\rm slm}}_{T}$ $\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T});\underline{\bm{v}}_{T}\bigr)\coloneqq sd^{{\rm slm}}_{T,{\rm!bc}}((\underline{\bm{u}}_{T},\underline{p}_{T});\underline{\bm{v}}_{T})$
${+}\int_{\partial T^{\rm N}}\bigl(\bm{u}_{\partial T}{\cdot}\bm{n}_{\partial T}\bigr)^{\oplus}\bigl(\rho_{\partial T}\bm{u}_{\partial T}{\cdot}{\bm{\pi}_{\partial T}^{k}}\bm{v}_{\partial T}\bigr){+}\int_{\partial T^{\rm N}}\bigl(\bm{u}_{\partial T}{\cdot}\bm{n}_{\partial T}\bigr)^{\ominus}\bigl(g_{\rm In}\bm{u}_{\partial T}{\cdot}{\bm{\pi}_{\partial T}^{k}}\bm{v}_{\partial T}\bigr)$
$-\int_{\partial T^{\rm N}}p_{\partial T}(\bm{v}_{\partial T}\cdot\bm{n}_{\partial T})+\int_{\partial T^{\rm N}}\bm{g}_{\rm N}\cdot\bm{v}_{\partial T}$
$+\int_{\partial T^{\rm D}}\bigl(\bm{g}_{\rm D}\cdot\bm{n}_{\partial T}\bigr)^{\oplus}\;(\rho_{\partial T}\bm{u}_{\partial T}\cdot\bm{v}_{\partial T})+\int_{\partial T^{\rm D}}\bigl(\bm{g}_{\rm D}\cdot\bm{n}_{\partial T}\bigr)^{\ominus}\;(g_{\rm In}\bm{g}_{\rm D}\cdot\bm{v}_{\partial T})$
$+\int_{\partial T^{\rm D}}\bigl((\bm{u}_{\partial T}-\bm{g}_{\rm D})\otimes\bm{n}_{\partial T}\bigr):\left(\nu{\mathcal{G}}_{T}\underline{\bm{v}}_{T}+\frac{\nu}{h_{F}}\bm{v}_{\partial T}\otimes\bm{n}_{\partial T}\right)+$
$-\int_{\partial T^{\rm D}}\nu{\mathcal{G}}_{T}\underline{\bm{u}}_{T}:(\bm{v}_{\partial T}\otimes\bm{n}_{\partial T}),$ (8a)
$sd^{\rm sms}_{T}$ $\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T});\underline{z}_{T}\bigr)\coloneqq sd^{\rm sms}_{T,{\rm!bc}}(\underline{\bm{u}}_{T},\underline{\rho}_{T};\underline{z}_{T})$
$+\int_{\partial T^{\rm N}}\bigl(\bm{u}_{\partial T}\cdot\bm{n}_{\partial T}\bigr)^{\oplus}\bigl(\rho_{\partial T}{z}_{\partial T}\bigr)+\int_{\partial T^{\rm N}}\bigl(\bm{u}_{\partial T}\cdot\bm{n}_{\partial T}\bigr)^{\ominus}\bigl(g_{\rm In}{z}_{\partial T}\bigr)$
$+\int_{\partial T^{\rm D}}\bigl(\bm{g}_{\rm D}\cdot\bm{n}_{\partial T}\bigr)^{\oplus}\bigl(\rho_{\partial T}{z}_{\partial T}\bigr)+\int_{\partial T^{\rm D}}\bigl(\bm{g}_{\rm D}\cdot\bm{n}_{\partial T}\bigr)^{\ominus}\bigl(g_{\rm In}{z}_{\partial T}\bigr),$ (8b)
$sd^{{\rm vol}}_{T}$ $(\underline{\bm{u}}_{T};\underline{q}_{T})\coloneqq sd^{{\rm vol}}_{T,{\rm!bc}}(\underline{\bm{u}}_{T};\underline{q}_{T})-\int_{\partial T^{\rm N}}(\bm{u}_{\partial T}\cdot\bm{n}_{\partial T})\,{q}_{\partial T}-\int_{\partial T^{\rm D}}(\bm{g}_{\rm D}\cdot\bm{n}_{\partial T})\,{q}_{\partial T}.$ (8c)
We remark that $\bm{u}_{\partial T}\cdot\bm{n}_{\partial T}=\bm{u}_{T}\cdot\bm{n}_{\partial T}$ on $\partial T_{\rm N}$, see Section 3.5 for details.
Given $(\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T})\in\underline{\bm{V}}_{T}^{k}\times\underline{Z}_{T}^{k}\times\underline{Q}_{T}^{k+1}$, the local spatial discretization of the unsteady momentum and mass conservation are such that, for all $\underline{\bm{v}}_{T}\in\underline{\bm{V}}_{T}^{k}$ and all $\underline{z}_{T}\in\underline{Z}_{T}^{k}$
$sd_{T}^{{\rm mnt}}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T});\underline{\bm{v}}_{T}\bigr)\coloneqq\int_{T}\frac{\partial({\rho}_{T}\bm{u}_{T})}{\partial t}\cdot\bm{v}_{T}+sd^{{\rm slm}}_{T}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T});\underline{\bm{v}}_{T}\bigr),$
$sd_{T}^{\rm mss}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T});\underline{z}_{T}\bigr)\coloneqq\int_{T}\frac{\partial\rho_{T}}{\partial t}z_{T}+h_{T}\int_{\partial T}\left(\frac{\partial\rho_{\partial T}}{\partial t}-\frac{\partial{\rho}_{T}}{\partial t}\right)({z}_{\partial T}-{z}_{T})+sd^{\rm sms}_{T}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T});\underline{z}_{T}\bigr).$ (9)
We remark that the second term of (9) is a time derivative stabilization designed to guarantee the robustness of the formulation whenever the normal velocity component happens to be null with respect to a mesh face. In this event, thanks to the additional stabilizations, the face density definition falls back to the average of the densities of the two elements sharing the face. We remark that a similar strategy was proposed to stabilize the linear momentum equations in the limit of vanishing viscosity, see [7] and [4].
The global residuals of the spatial discretization $r^{\rm mnt}_{h}((\underline{\bm{u}}_{h},\underline{\rho}_{h},\underline{p}_{h});\cdot):\underline{\bm{V}}^{k}_{h}\to\mathbb{R}$, $r^{\rm mss}_{h}((\underline{\bm{u}}_{h},\underline{\rho}_{h});\cdot):\underline{Z}^{k}_{h}\to\mathbb{R}$, and $r^{\rm vol}_{h}(\underline{\bm{u}}_{h};\cdot):\underline{Q}^{k+1}_{h}\to\mathbb{R}$ are obtained by cell-by-cell assembly of the local discretizations, i.e.: For all $\underline{\bm{v}}_{h}\in\underline{\bm{V}}_{h}^{k}$, all $\underline{z}_{h}\in\underline{Z}_{h}^{k}$ and all $\underline{q}_{h}\in\underline{Q}_{h}^{k+1}$,
$r_{h}^{\rm mnt}\bigl((\underline{\bm{u}}_{h},\underline{\rho}_{h},\underline{p}_{h});\underline{\bm{v}}_{h}\bigr)$ $\coloneqq\sum_{T\in\mathcal{T}_{h}}sd^{\rm mnt}_{T}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T});\underline{\bm{v}}_{T}\bigr),$ (10a)
$r^{\rm mss}_{h}\bigl((\underline{\bm{u}}_{h},\underline{\rho}_{h});\underline{z}_{h}\bigr)$ $\coloneqq\sum_{T\in\mathcal{T}_{h}}sd^{\rm mss}_{T}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T});\underline{z}_{T}\bigr),$ (10b)
$r_{h}^{\rm vol}(\underline{\bm{u}}_{h};\underline{q}_{h})$ $\coloneqq\sum_{T\in\mathcal{T}_{h}}sd^{\rm vol}_{T}(\underline{\bm{u}}_{T};\underline{q}_{T}).$ (10c)
The global problem reads: Find $(\underline{\bm{u}}_{h},\underline{\rho}_{h},\underline{p}_{h})\in\underline{\bm{V}}_{h}^{k}\times\underline{Z}_{h}^{k}\times\underline{Q}^{k+1}_{h}$ such that
$r_{h}^{\rm mnt}\bigl((\underline{\bm{u}}_{h},\underline{\rho}_{h},\underline{p}_{h});\underline{\bm{v}}_{h}\bigr)$ $=0$ $\forall\underline{\bm{v}}_{h}\in\underline{\bm{V}}_{h}^{k},$ (11)
$r_{h}^{\rm mss}\bigl((\underline{\bm{u}}_{h},\underline{\rho}_{h});\underline{z}_{h}\bigr)$ $=0$ $\forall\underline{z}_{h}\in\underline{Z}_{h}^{k},$
$r_{h}^{\rm vol}(\underline{\bm{u}}_{h};\underline{q}_{h})$ $=0$ $\forall\underline{q}_{h}\in\underline{Q}_{h}^{k+1}.$
3.5 H(div)-conformity and independence from irrotational body forces
Following [8] we consider the HHO formulation of the volume conservation equation. The requirement $r_{h}^{\rm vol}(\underline{\bm{u}}_{h};\underline{q}_{h})=0$ for each $\underline{\bm{u}}_{h}\in\underline{\bm{V}}_{h}$ and for all $\underline{q}_{h}\in\underline{Q}^{k+1}_{h}$ reads
$0=\sum_{T\in\mathcal{T}_{h}}\left(-\int_{T}{q}_{T}\;\nabla\cdot\bm{u}_{T}+\int_{\partial T}{q}_{\partial T}\bm{u}_{T}\cdot\bm{n}_{TF}-\int_{\partial T^{\rm N}}{q}_{\partial T}\bm{u}_{\partial T}\cdot\bm{n}_{\partial T}-\int_{\partial T^{\rm D}}{q}_{\partial T}\bm{g}_{\rm D}\cdot\bm{n}_{\partial T}\right).$
Among all the $\underline{q}_{h}\in\underline{Q}^{k+1}_{h}$, we consider the following relevant choices
  1. 1.
    $\underline{q}_{h}=\left((q_{T})_{T\in\mathcal{T}_{h}},(0)_{F\in\mathcal{F}_{h}}\right)$, accordingly, $\forall T\in\mathcal{T}_{h}$ and $\forall{q}_{T}\in\mathbb{P}_{d}^{k}(T)$, we get $\int_{T}{q}_{T}\;\nabla\cdot\bm{u}_{T}=0,\;$ i.e., $\nabla\cdot\bm{u}_{T}=0,\forall\bm{x}\in T$, since $\nabla\cdot\bm{u}_{T}\in\mathbb{P}_{d}^{k}(T)$;
  2. 2.
    $\underline{q}_{h}=\left((1)_{T\in\mathcal{T}_{h}},(0)_{F\in\mathcal{F}_{h}}\right)$, accordingly, after integration by parts, $\forall T\in\mathcal{T}_{h}$, we get $\int_{\partial T}\bm{u}_{T}\cdot\bm{n}_{TF}=0;$
  3. 3.
    $\underline{q}_{h}=\left((0)_{T\in\mathcal{T}_{h}},(q_{F})_{F\in\mathcal{F}_{h}}\right)$, accordingly, $\forall{q}_{F}\in\mathcal{P}^{k+1}(F)$, we get
    $\forall F\in\mathcal{F}_{h}^{\rm D},$ $\int_{F}{q}_{F}(\bm{u}_{T}-\bm{g}_{\rm D}){\cdot}\bm{n}_{TF}=0,$ $\;\text{i.e., }\;\bm{u}_{T}{\cdot}\bm{n}_{TF}{=}\pi_{F}^{k+1}(\bm{g}_{D}{\cdot}\bm{n}_{TF}),$
    $\forall F\in\mathcal{F}_{h}^{\rm N},$ $\int_{F}{q}_{F}(\bm{u}_{T}-\bm{u}_{F}){\cdot}\bm{n}_{TF}=0,$ $\;\text{i.e., }\;\bm{u}_{T}{\cdot}\bm{n}_{TF}{=}\bm{u}_{F}{\cdot}\bm{n}_{TF},$
    $\forall F\in\mathcal{F}_{h}^{\rm i},$ $\int_{F}{q}_{F}\;(\bm{u}_{T}-\bm{u}_{T^{\prime}}){\cdot}\bm{n}_{TF}=0,$ $\;\text{i.e., }\;\bm{u}_{T}{\cdot}\bm{n}_{TF}{=}\bm{u}_{T^{\prime}}{\cdot}\bm{n}_{TF},$
    where the equalities involving normal velocity components hold true $\forall\bm{x}\in F$. Note that a Neumann face is such that $\bm{u}_{F}\in\mathcal{P}^{k+1}(F)$. Note that an internal face is such that $F=\partial T\cap\partial T^{\prime}$, with $^{\prime}\in\mathcal{T}_{h}, TTT\neq T^{\prime}$ and $\bm{n}_{TF}=-\bm{n}_{T^{\prime}F}$.
The above procedure allows to conclude that the velocity field is pointwise divergence-free and $\bm{H}(\operatorname{div})$-conforming, and to infer that conservation of volume is enforced cell-by-cell.
[…]
$\int_{T}\bm{f}\cdot\bm{v}_{T}$ $=\int_{T}\bm{\ell}\cdot\bm{v}_{T}+\int_{T}(\nabla\psi)\cdot\bm{v}_{T}$ (12)
$=\int_{T}\bm{\ell}\cdot\bm{v}_{T}-\int_{T}\psi\;(\nabla\cdot\bm{v}_{T})+\int_{\partial T}\psi\;\bm{v}_{T}\cdot\bm{n}_{\partial T}$
$=\int_{T}\bm{\ell}\cdot\bm{v}_{T}-\int_{T}\pi^{k}_{T}\psi\;(\nabla\cdot\bm{v}_{T})+\int_{\partial T}\pi^{k+1}_{\partial T}\psi\;\bm{v}_{T}\cdot\bm{n}_{\partial T}$
$=\int_{T}\bm{\ell}\cdot\bm{v}_{T}+\int_{T}{\bm{\mathfrak{g}}}_{T}^{k+1}(\underline{I}_{{Q},T}\psi)\cdot\bm{v}_{T},\quad$
where, for all $q\in H^{1}(\Omega)$, $\underline{I}_{\underline{Q},T}q\coloneqq\bigl(\pi_{T}^{k}q,\pi_{\partial T}^{k+1}q\bigr)$, is the local interpolation operator onto $\underline{Q}_{T}$. Note that integration by parts allows to pass from the first to the second line. Moreover, since $\bm{v}_{T}\in\mathcal{P}^{k+1}(T)$, $\nabla\cdot\bm{v}_{T}\in\mathcal{P}^{k}(T)$, and $\bm{n}_{TF}$ is a constant vector over $F$, $L^{2}$-projectors can be introduced on the third line.
Introducing (12) in $sd^{{\rm slm}}_{T,{\rm!bc}}$, see in particular (5), it becomes clear that the irrotational part of the body force can be combined with the pressure gradient discretization to obtain an equivalent reformulation of the momentum equation spatial dicretization. The procedure required to achieve the desired reformulation can be summarized as follows: replace
  1. 1.
    $\int_{T}{\bm{\mathfrak{g}}}_{T}^{k+1}\underline{p}_{T}\cdot\bm{v}_{T}$ with $\int_{T}{\bm{\mathfrak{g}}}_{T}^{k+1}(\underline{p}_{T}-\underline{I}_{{Q},T}\psi)\cdot\bm{v}_{T}$
  2. 2.
    $\int_{T}\bm{f}\cdot\bm{v}_{T}$ with $\int_{T}\bm{\ell}\cdot\bm{v}_{T}$,
  3. 3.
    -$\int_{\partial T^{\rm N}}p_{\partial T}\bm{v}_{\partial T}\cdot\bm{n}_{\partial T}$ with $\int_{\partial T^{\rm N}}(p_{\partial T}-\pi_{\partial T}^{k+1}\psi)\bm{v}_{\partial T}\cdot\bm{n}_{\partial T}$
  4. 4.
    $\int_{\partial T^{\rm N}}\bm{g}_{\rm N}\cdot\bm{v}_{\partial T}$ with $\int_{\partial T^{\rm N}}(\bm{g}_{\rm N}-\pi_{\partial T}^{k+1}\psi\bm{n}_{\partial T})\cdot\bm{v}_{\partial T}$
in the definition of $sd^{{\rm slm}}_{T}$, see (8a) and (5). Note that the last two occurrences in the list above are required to make Neumann boundary conditions consistent with the reformulation and can be interpreted as adding and subtracting $\int_{\partial T^{\rm N}}\pi_{F}^{k+1}\psi\,\bm{v}_{\partial T}\cdot\bm{n}_{\partial T}$ to $sd^{{\rm slm}}_{T}$. Interestingly, the refactored HHO discretization, can be rewritten in terms of the modified pressure unknown $\widetilde{\underline{p}}_{h}\coloneqq\bigl(({p}_{T}-\pi_{T}^{k}\psi)_{T\in\mathcal{T}_{h}},({p}_{F}-\pi_{F}^{k+1}\psi)_{F\in\mathcal{F}_{h}}\bigr)$, being otherwise indistinguishable from the original version. This procedure allows to conclude that only the pressure field is affected by the irrotational part of body forces.
3.6 ESDIRK temporal discretization
The temporal discretization must take into account that the HHO space discretization of the variable density problem is formulated in terms of primitive variables $\bm{u},\rho$ rather then conservative variables $\rho\bm{u},\rho$. This latter choice allows to obtain a $\bm{H}(\operatorname{div})$-conforming velocity field and achieve pressure-robustness. In order to rewrite the variable density problem as a system of Ordinary Differential Equations (ODEs), we first introduce the algebraic problem corresponding to the semi-discrete HHO formulation. For the sake of notation, in this section, we disregard the time derivative stabilization in (9).
The unknowns for a mesh cell $T\in\mathcal{T}_{h}$ correspond to the coefficients of the expansions of the velocity, density and pressure in the selected local bases. Assuming that the unknowns are ordered so that cell velocities come first and boundary velocities next, these coefficients are collected in the vectors
$\underline{U}_{T}=\begin{bmatrix}{\sf U}_{T}\\ {\sf U}_{\partial T}\end{bmatrix}$, $\underline{R}_{T}=\begin{bmatrix}{\sf R}_{T}\\ {\sf R}_{\partial T}\end{bmatrix}$, and $\underline{P}_{T}=\begin{bmatrix}{\sf P}_{T}\\ {\sf P}_{\partial T}\end{bmatrix}$,
where the block partition is the one naturally induced by the selected ordering of unknowns. For convenience we collect element and face unknowns in the vectors
${{\sf W}}_{T}=\begin{bmatrix}{\sf U}_{T}\\ {\sf R}_{T}\\ {\sf P}_{T}\end{bmatrix}$, ${{\sf W}}_{T}^{!p}=\begin{bmatrix}{\sf U}_{T}\\ {\sf R}_{T}\\ {\sf 0}\end{bmatrix}$, ${{\sf W}}_{\partial T}=\begin{bmatrix}{\sf U}_{\partial T}\\ {\sf R}_{\partial T}\\ {\sf P}_{\partial T}\end{bmatrix}$,  and we define $\underline{W}_{T}\coloneqq\begin{bmatrix}{\sf W}_{T}\\ {\sf W}_{\partial T}\end{bmatrix}$.
Notice that the vector of element unknowns omitting the pressure in ${{\sf W}}_{T}^{!p}$ have been introduced to conveniently express time derivatives in what follows. We also introduce the notation $\underline{W}_{\mathcal{T}_{h}}$ for the global vector collecting all element and face unknowns.
Let’s introduce the vector representations
$\underline{D}_{T}^{\rm slm}=\begin{bmatrix}{\sf D}_{T}^{\rm slm}\\ {\sf D}_{\partial T}^{\rm slm}\end{bmatrix},\;\underline{D}_{T}^{\rm sms}=\begin{bmatrix}{\sf D}_{T}^{\rm sms}\\ {\sf D}_{\partial T}^{\rm sms}\end{bmatrix},\;\underline{D}_{T}^{\rm vol}=\begin{bmatrix}{\sf D}_{T}^{\rm vol}\\ {\sf D}_{\partial T}^{\rm vol}\end{bmatrix}$,
of the local spatial discretizations $sd^{\rm slm}_{T},\,sd^{\rm sms}_{T},\,sd^{\rm vol}_{T}$ defined in (8). The block partition is the one induced by mimicking the selected ordering of unknowns for the expansion bases. For convenience we collect element and face components in the vectors
${\sf D}_{T}^{s}=\begin{bmatrix}{\sf D}_{T}^{\rm slm}\\ {\sf D}_{T}^{\rm sms}\\ {\sf D}_{T}^{\rm vol}\end{bmatrix}$,   ${\sf D}_{\partial T}^{s}=\begin{bmatrix}{\sf D}_{\partial T}^{\rm slm}\\ {\sf D}_{\partial T}^{\rm sms}\\ {\sf D}_{\partial T}^{\rm vol}\end{bmatrix}$,  and we define $\underline{D}_{T}^{s}=\begin{bmatrix}{\sf D}_{T}^{s}\\ {\sf D}_{\partial T}^{s}\end{bmatrix}$.
The algebraic counterpart of problem (11) reads: find $\underline{W}_{\mathcal{T}_{h}}$ such that
$\sum_{T\in\mathcal{T}_{h}}\bm{M}_{T}\frac{d{\sf W}_{T}^{!p}}{dt}+\sum_{T\in\mathcal{T}_{h}}{\sf D}_{T}^{s}(\underline{W}_{T})={\sf 0},\quad\text{and}\quad\sum_{T\in\mathcal{T}_{h}}{\sf D}_{\partial T}^{s}(\underline{W}_{T})={\sf 0},$ (13)
where
$\bm{M}_{T}({\sf W}_{T})=\begin{bmatrix}\int_{T}{\rho}_{T}\;\widetilde{\sf{v}}_{T}\otimes\widetilde{\sf{v}}_{T}&\int_{T}\bm{u}_{T}\;\widetilde{\sf{v}}_{T}\otimes\widetilde{\sf{z}}_{T}&\bm{0}\\ \bm{0}&\int_{T}\widetilde{\sf{z}}_{T}\otimes\widetilde{\sf{z}}_{T}&\bm{0}\\ \bm{0}&\bm{0}&\bm{I}\\ \end{bmatrix}$,
and $\widetilde{\sf{v}}_{T}$, $\widetilde{\sf{z}}_{T}$, $\widetilde{\sf{q}}_{T}$ are the vectors collecting local cell bases, i.e.,
$\begin{array}[]{c}\operatorname{span}\left\{\widetilde{\sf{v}}^{T}\right\}=\mathcal{P}^{k+1}(T)^{d}\\ \operatorname{span}\left\{\widetilde{\sf{z}}_{T}\right\}=\mathcal{P}^{k}(T)\\ \operatorname{span}\left\{\widetilde{\sf{q}}_{T}\right\}=\mathcal{P}^{k}(T)\\ \end{array}\quad\text{and, for any $\bm{x}\in T$}\quad\begin{array}[]{c}\bm{u}_{T}|_{\bm{x}}={\sf U}_{T}\cdot\widetilde{\sf{v}}_{T}|_{\bm{x}}\\ {\rho}_{T}|_{\bm{x}}={\sf R}_{T}\cdot\widetilde{\sf{z}}_{T}|_{\bm{x}}\\ {p}_{T}|_{\bm{x}}={\sf P}_{T}\cdot\widetilde{\sf{q}}_{T}|_{\bm{x}}\\ \end{array}.$
Problem (13) can be finally rewritten as a system of ODEs: find $\underline{W}_{\mathcal{T}_{h}}$ such that
$\sum_{T\in\mathcal{T}_{h}}\frac{d{\sf W}_{T}^{!p}}{dt}+\sum_{T\in\mathcal{T}_{h}}\bm{M}_{T}^{-1}{\sf D}^{s}_{T}(\underline{W}_{T})={\sf 0},\quad\text{and}\quad\sum_{T\in\mathcal{T}_{h}}{\sf D}_{\partial T}^{s}(\underline{W}_{T})={\sf 0},$ (14)
where, $\bm{M}_{T}^{-1}({\sf W}_{T}){=}$
$\begin{bmatrix}\left(\int_{T}{\rho}_{T}\widetilde{\sf{v}}_{T}\otimes\widetilde{\sf{v}}_{T}\right)^{-1}&-\left(\int_{T}{\rho}_{T}\widetilde{\sf{v}}_{T}\otimes\widetilde{\sf{v}}_{T}\right)^{-1}\left(\int_{T}\bm{u}_{T}\widetilde{\sf{v}}_{T}\otimes\widetilde{\sf{z}}_{T}\right)\left(\int_{T}\widetilde{\sf{z}}_{T}\otimes\widetilde{\sf{z}}_{T}\right)^{-1}&\bm{0}\\ \bm{0}&(\int_{T}\widetilde{\sf{z}}_{T}\otimes\widetilde{\sf{z}}_{T})^{-1}&\bm{0}\\ \bm{0}&\bm{0}&\bm{I}\\ \end{bmatrix}$.
Based on multistage ESDIRK time marching strategies, the solutions $\underline{W}_{\mathcal{T}_{h}}|_{t+\delta t}$ are computed considering an approximated time integration formula for the global residuals in (14) over the time interval $\delta t$:
$\sum_{T\in\mathcal{T}_{h}}\frac{\delta{\sf W}_{T}^{!p}}{\delta t}+\sum_{T\in\mathcal{T}_{h}}\sum_{i=1}^{s}b_{i}\,\bm{M}_{T}^{-1}({\sf W}_{T}|_{t+\delta t_{i}}){\sf D}_{T}^{s}(\underline{W}_{T}|_{t+\delta t_{i}})$ $={\sf 0},$ (15)
$\sum_{T\in\mathcal{T}_{h}}\sum_{i=1}^{s}b_{i}\,{\sf D}_{\partial T}^{s}(\underline{W}_{T}|_{t+\delta t_{i}})$ $={\sf 0},$
where $\delta{\sf W}_{X}\coloneqq\underline{W}_{X}|_{t+\delta t}-\underline{W}_{X}|_{t}$, and, for all $T\in\mathcal{T}_{h}$, $\underline{W}_{T}|_{t+\delta t_{i}}$ are the solutions of the ESDIRK $i$th-stage, with $\delta t_{i}=\delta t\;\sum_{j=1}^{i}a_{ij}$. The number of stages $s$ and the real coefficients $a_{ij}$, with $i=1,\dots,s$ and $j=1,\dots,i$, uniquely define an ESDIRK formulation. In particular, for all schemes belonging to the ESDIRK family, it holds that $a_{11}=0$ and $b_{i}=a_{si}$, for $i=1,\dots,s$. As a consequence, since $\delta t_{1}=0$ and $\delta t_{s}=\delta t$, the approximated time integration formula in (15) is equivalent to the last ESDIRK stage and $\underline{W}_{\mathcal{T}_{h}}|_{t+\delta t}$ is the last stage solution. The stages solutions are built incrementally, starting from the first stage, moving to second, and so on up to the last stage. To complete the presentation of the time marching strategy we need to specify how the ESDIRK $i$th-stage solutions are computed.
Given $\underline{W}_{\mathcal{T}_{h}}|_{t+\delta t_{j}}$, with $j=1,\dots,i$, the space-time discretization of the ESDIRK $i$th-stage reads
${\sf D}_{i,T}^{st}(\underline{W}_{T}(t))$ $\coloneqq\frac{\delta_{i}{\sf W}_{T}^{!p}}{\delta t}+\sum_{j=1}^{i}a_{ij}\,\bm{M}_{T}^{-1}({\sf W}_{T}|_{t+\delta t_{j}}){\sf D}_{T}^{s}(\underline{W}_{T}|_{t+\delta t_{j}}),$ (16a)
${\sf D}_{i,\partial T}^{st}(\underline{W}_{T}(t))$ $\coloneqq\sum_{j=1}^{i}a_{ij}\,{\sf D}_{\partial T}^{s}(\underline{W}_{T}|_{t+\delta t_{j}}),$ (16b)
where $\delta_{i}{\sf W}_{X}\coloneqq{\sf W}_{X}|_{t+\delta t_{i}}-{\sf W}_{X}|_{t}$. We remark that, the spatial discretizations $\underline{D}_{T}^{s}(\underline{W}_{T}(t))$ in (15)-(16) are written omitting the dependence from the boundary conditions and forcing terms. In doing so, we implicitly assume that $\bm{f}$, $\bm{g}_{\rm N},\bm{g}_{\rm D}$ and $g_{\rm In}$ are evaluated at the same time points as the variables $\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T}$.
The ESDIRK $i$th-stage problem, that is the time discrete counterpart of problem (11), reads: Given $\underline{W}_{\mathcal{T}_{h}}|_{t+\delta t_{j}}$, with $j=1,\dots,i-1$, find $\underline{W}_{\mathcal{T}_{h}}|_{t+\delta t_{i}}$ such that
$\sum_{T\in\mathcal{T}_{h}}{\sf D}_{i,T}^{st}(\underline{W}_{T}(t))={\sf 0}\quad\text{and}\quad\sum_{T\in\mathcal{T}_{h}}{\sf D}_{i,\partial T}^{st}(\underline{W}_{T}(t))={\sf 0}.\\$ (17)
where the left hand sides are the global element and face residuals, respectively, of the ESDIRK $i$th-stage. The solution of the above problem can be sought by means of Newton’s method, as detailed in the next section.
3.7 Algebraic expression for Newton’s method and static condensation
In order to streamline the computation of the Jacobian matrix we scale the ESDIRK residual equations in (17) by $\frac{1}{a_{ii}}$, to obtain the following equivalent reformulation of the $i$-th stage space-time discretization
${\sf\widehat{D}}_{i,T}^{st}(\underline{W}_{T}(t))$ $\coloneqq\frac{\delta_{i}{\sf W}_{T}^{!p}}{a_{ii}\,\delta t}+\bm{M}_{T}^{-1}({\sf W}_{T}|_{t+\delta t_{i}}){\sf D}_{T}^{s}(\underline{W}_{T}|_{t+\delta t_{i}})p$
$\qquad\qquad\qquad\qquad+\sum_{j=1}^{i-1}\frac{a_{ij}}{a_{ii}}\,\bm{M}_{T}^{-1}({\sf W}_{T}|_{t+\delta t_{j}}){\sf D}_{T}^{s}(\underline{W}_{T}|_{t+\delta t_{j}})$ (18a)
${\sf\widehat{D}}_{i,\partial T}^{st}(\underline{W}_{T}(t))$ $\coloneqq{\sf D}_{\partial T}^{s}(\underline{W}_{T}|_{t+\delta t_{i}})+\sum_{j=1}^{i-1}\frac{a_{ij}}{a_{ii}}\,{\sf D}_{\partial T}^{s}(\underline{W}_{T}|_{t+\delta t_{j}})$ (18b)
Newton’s method for solving problem (17) reads: Find $\delta_{i}\underline{W}_{\mathcal{T}_{h}}$ such that
[…]
where $\bm{\widehat{J}}_{XY}\coloneqq\frac{\partial{\sf\widehat{D}}^{st}_{i,X}(\underline{W}_{T}(t))}{\partial{\sf W}_{Y}|_{t+\delta t_{i}}}$ are the blocks of the local Jacobian matrix and $\bm{M}_{i,T}\coloneqq\bm{M}_{T}({\sf W}_{T}|_{t+\delta t_{i}})$, replace $\underline{W}_{\mathcal{T}_{h}}|_{t+\delta t_{i}}\leftarrow\underline{W}_{\mathcal{T}_{h}}|_{t+\delta t_{i}}+\delta_{i}\underline{W}_{\mathcal{T}_{h}}$ until $\delta_{i}\underline{W}_{\mathcal{T}_{h}}$ is small enough. Note that the each of the element residual equations in (19) has been multiplied by $\bm{M}_{i,T}$ in order to further simplify the code implementation of the Jacobian matrix. In particular, since
$\frac{\partial\bm{M}_{i,T}^{-1}}{\partial{\sf W}_{T}|_{t+\delta t_{i}}}=-\bm{M}_{i,T}^{-1}\frac{\partial\bm{M}_{i,T}}{\partial{\sf W}_{T}|_{t+\delta t_{i}}}\bm{M}_{i,T}^{-1},\quad\text{and}\quad\frac{\partial\bm{M}_{i,T}^{-1}}{\partial{\sf W}_{\partial T}|_{t+\delta t_{i}}}=\bm{0},$
the equations systems in (19) can be rewritten and follows
[…]
where
$\bm{M}_{i,T}\bm{\widehat{J}}_{TT}=\frac{\bm{M}_{i,T}}{a_{ii}\,\delta t_{i}}\frac{\partial{\sf W}_{T}^{!p}|_{t+\delta t_{i}}}{\partial{\sf W}_{T}|_{t+\delta t_{i}}}-\frac{\partial\bm{M}_{i,T}}{\partial{\sf W}_{T}|_{t+\delta t_{i}}}\bm{M}_{i,T}^{-1}{\sf D}^{s}_{T}(\underline{W}_{T}|_{t+\delta t_{i}})+\bm{J}_{TT}$
and $\bm{J}_{XY}=\frac{\partial{\sf D}^{s}_{X}(\underline{W}_{T}|_{t+\delta t_{i}})}{\partial{\sf W}_{Y}|_{t+\delta t_{i}}}$. Notice that the spatial discretizations at the $j$th-stages, j=1,,i1j=1,...,i{-}1, that is the last terms in (18a) and (18b), do not appear in the Jacobian matrix related to the $i$th-stage.
Static condensation is applied to the equations systems in (20) in a cell-by-cell fashion. The only unknowns are globally coupled are those collected in the subvector $\delta_{i}{\sf W}_{\partial T}$ while the remaining unknowns collected in $\delta_{i}{\sf W}_{T}$ can be eliminated by expressing them in terms of the former. After performing this local elimination, each Newton method iteration amounts at solving
$\sum_{T\in\mathcal{T}_{h}}\left({\sf\widehat{D}}^{st}_{i,\partial T}{-}\bm{J}_{\partial TT}\bigl(\bm{M}_{i,T}\bm{\widehat{J}}_{TT}\bigr)^{-1}\bigl(\bm{M}_{i,T}{\sf\widehat{D}}^{st}_{i,T}\bigr)\\ \right){=}\sum_{T\in\mathcal{T}_{h}}\bigl(\bm{S}_{T}\delta_{i}{\sf W}_{\partial T}\\ \bigr),$
where $\bm{S}_{T}$ denotes the Schur complement of the top left block of the left-hand side matrix in (20), that is,
$\bm{S}_{T}\coloneqq\bm{J}_{\partial T\partial T}-\bm{J}_{\partial TT}\bigl(\bm{M}_{i,T}\bm{\widehat{J}}_{TT}\bigr)^{-1}\bm{J}_{T\partial T}$
When increasing the polynomial degree $k$, the dimension of polynomial spaces over mesh cells grows much faster that the dimension of polynomial spaces over mesh faces. Accordingly, the possibility to statically condense the cells unknowns is the key for achieving efficiency when high-order accurate formulation are employed.
4 Numerical results
Let us briefly describe the numerical setup common to the test cases presented in what follows. Numerical integration over mesh cells and mesh faces is performed over standardized reference cells based on Gaussian quadrature rules. Reference-to-physical-frame mappings from reference entities to mesh entities are defined by means of Lagrange polynomials. Discrete polynomial spaces over mesh cells and mesh faces are spanned by orthonormal modal bases. Specifically, we rely on physical frame polynomial spaces over mesh cells and reference frame polynomial spaces over mesh faces, as proposed in [6]. Orthogonalization in the physical frame is performed cell-by-cell starting from a monomial basis defined in a local reference frame aligned with the principal axes of inertia of each mesh cell, as described in [1]. Note that, on simplicial meshes, since planar faces and affine mappings go hand-in-hand, reference and physical frame polynomials provide the same accuracy, see [5] for details.
We remark that Newton’s method is solved up to machine precision in order to guarantee exact conservation of volume at the discrete level, accordingly the linear solver stopping criterion does not influence the numerical results. In order to further improve the performance of the iterative linear solver required for the solution of each Newton step we rely on the $p$-multilevel preconditioner strategy proposed in [9]. In particular, we apply one sweep of $p$-multigrid V-cycle as a preconditioner for the FGMRES (flexible GMRES) iteration. As a smoothing strategy within the multigrid V-cycle preconditioner we rely on a few iterations of GMRES. All applications of prolongation and restriction operators involved in the V-cycle iteration are performed matrix-free, that is, without assembling the global sparse matrices associated to the operators, see [9] for details.
4.1 Kovasznay test case
As a first validation step we seek to numerically demonstrate pure advection of the density by takling the Kovasznay flow problem [18]. The analytical solution of the incompressible Navier–Stoke equations, which describes steady flow behing a grid made of equally spaces parallel rods for a fluid with constant density, reads as follows
$\bm{u}$ $=\left[1-\mathrm{e}^{\kappa x}\cos{\left(2\pi y\right)}\right]\bm{i}+\dfrac{\kappa}{2\pi}\mathrm{e}^{\kappa x}\sin{\left(2\pi y\right)}\bm{j},$ (21)
$p$ $=p_{0}-\dfrac{1}{2}\mathrm{e}^{2\kappa x},$
$\rho$ $=1,$
where $p_{0}\in\mathbb{R}$ is an arbitrary constant and the parameter $\kappa$ depends on the Reynolds number
$\kappa=\dfrac{\mathrm{Re}}{2}-\sqrt{\dfrac{\mathrm{Re}^{2}}{4}+4\pi^{2}}.$
The Kovasznay flow problem is defined on the bi-unit square computational domain $\Omega=\left(-0.5,1.5\right){\times}\left(0,2\right)$ and boundary conditions are derived from the Kovasznay analytical solution (21) at $\mathrm{Re}{=}\frac{1}{\mu}{=}40$. In particular, a Neumann boundary condition is imposed at the outflow boundary (right side of the square domain) while Dirichlet boundary conditions are imposed on the remaining boundaries. Since the density is uniform and purely advected, we expect the density solution at the steady state to match, up to numerical precision, the unit density field employed as initial guess.
We apply the variable density HHO formulation to a $h$-refined mesh sequence composed of regular triangular elements and compute numerical solutions for several polynomial degrees $k=0,1,2,3,4$. Triangular elements meshes are generated by means of the Delaunay algorithm and, at each step of the mesh sequence, a four fold increase of the mesh cardinality is approximately targeted. Errors in $L^{2}$-norm and $h$-convergence rates are computed testing against the analytical solution and tabulated for the approximated velocity, velocity gradients and pressure fields, see table 1. A k+1k{+}1 rate of error reduction is observed for the pressure error in $L^{2}$-norm. Since the flow is diffusion dominated at this Reynolds number, the velocity and velocity gradient errors in the $L^{2}$-norm show rates of convergence of k+2k{+}2 and k+1k{+}1, respectively. The errors in $L^{2}$-norm for the velocity divergence, also reported table 1, show that, the incompressibility constraint is enforced up to machine precision. Moreover, as confirmed by the density errors in $L^{2}$-norm reported in table 1, we are able to assess that a unit density field is obtained at all the polynomial degrees and on all the computations grids, compatibly with the claim that density is purely advected.
$\operatorname{card}\mathcal{T}_{h}$ $\mathfrak{E}\,{u_{h}}$ rate $\mathfrak{E}\,{\nabla u_{h}}$ rate $\mathfrak{E}\,{p_{h}}$ rate $\mathfrak{E}\,{\rho_{h}}$ $\mathfrak{E}\,{\nabla{\cdot}u_{h}}$
$k=0$
50 0.456 4.828 0.217 6.6e-16 1.6e-16
192 0.123 1.95 2.853 0.78 0.0999 1.16 7.7e-16 3.3e-16
810 0.0291 2.00 1.390 1.00 0.0439 1.14 1.0e-15 2.4e-16
3126 0.00828 1.86 0.690 1.04 0.0218 1.03 1.1e-15 2.7e-16
12706 0.00207 1.97 0.340 1.01 0.0109 0.99 1.5e-15 2.7e-16
$k=1$
50 0.105 2.280 0.0536 3.1e-14 3.0e-16
192 0.0116 3.28 0.550 2.11 0.00629 3.19 8.5e-15 3.8e-16
810 0.00139 2.94 0.135 1.95 0.00114 2.36 3.2e-14 3.3e-16
3126 0.000197 2.90 0.0370 1.92 0.000286 2.06 3.3e-14 3.2e-16
12706 2.46e-05 2.97 0.00904 2.01 7.27e-05 1.95 2.2e-14 3.2e-16
$k=2$
50 0.0319 0.841 0.0186 5.0e-14 4.4e-16
192 0.00168 4.37 0.107 3.05 0.000739 4.80 7.2e-14 4.5e-16
810 8.75e-05 4.11 0.0124 3.00 5.51e-05 3.61 2.9e-14 3.8e-16
3126 5.65e-06 4.06 0.00161 3.02 6.36e-06 3.20 3.6e-14 4.0e-16
12706 3.56e-07 3.94 0.000204 2.95 7.84e-07 2.99 5.2e-14 3.9e-16
$k=3$
50 0.00285 0.136 0.00130 6.1e-14 4.7e-16
192 9.52e-05 5.05 0.00991 3.90 4.96e-05 4.86 1.0e-13 4.7e-16
810 3.14e-06 4.74 0.000663 3.76 2.10e-06 4.39 4.4e-14 3.6e-16
3126 1.23e-07 4.80 5.04e-05 3.81 1.34e-07 4.07 6.6e-14 4.0e-16
12706 3.90e-09 4.92 3.15e-06 3.95 8.22e-09 3.99 2.8e-13 3.8e-16
$k=4$
50 0.000739 0.0429 0.000349 1.8e-14 4.9e-16
192 9.45e-06 6.48 0.00123 5.27 4.36e-06 6.52 2.4e-14 5.0e-16
810 1.49e-07 5.77 3.87e-05 4.81 1.04e-07 5.18 5.3e-14 4.1e-16
3126 2.49e-09 6.06 1.32e-06 5.00 3.36e-09 5.09 1.3e-13 4.5e-16
12706 3.95e-11 5.91 4.18e-08 4.93 1.04e-10 4.95 2.0e-13 4.4e-16
Table 1: Kovasnzay test case: spatial convergence rates for $k=0,1,2,3,4$ HHO formulations over a Delaunay type $h$-refined triangular elements mesh sequence.
4.2 Guermond and Quartapelle manufactured solution
In order to validate spatial and temporal convergence rates for the HHO-ESDIRK formulation we consider the variable density manufactured solution proposed by Guermond and Quartapelle [16]. The analytical velocity, pressure and density fields read
$\bm{u}$ $=(-y\bm{i}+x\bm{j})\cos{\left(t\right)},$
$p$ $=\sin{\left(x\right)}\sin{\left(y\right)}\sin{\left(t\right)},$
$\rho$ $=2+x\cos{\left({\sin{\left(t\right)}}\right)}+y\sin{\left({\sin{\left(t\right)}}\right)}.$
and the forcing term $\bm{f}$ within the momentum equation is set according to the analytical solution. Dirichlet and Neumann boundary conditions are imposed on the boundary of the unit square domain $\Omega=\left(0,1\right)^{2}$, in particular two of the four edges are of Dirichlet type and the other two are of Neumann type. Time integration is carried out in the interval $(0,\frac{7}{4}\pi]$. Accordingly, the errors in $L^{2}$ norm for the velocity, the velocity gradients, the velocity divergence, the pressure and the density are computed based on the exact solution at the final time $t_{\rm F}=\frac{7}{4}\pi$. The velocity and density fields are initialized interpolating the exact solution. We rely on triangular cells meshes obtained subdividing in two right triangles the square cells of $N^{2}$ Cartesian grids, with $N$ cells per Cartesian direction. Accordingly, the mesh cardinality is $\operatorname{card}(\mathcal{T}_{h})=2\,N^{2}$ when $h=\frac{\sqrt{2}}{N}$.
The spatial convergence results are reported in Table 2 for $k=0,1,2$ HHO formulations considering five grids of a triangular mesh sequence with $N=4\times{2^{i}}$, $i=0,\dots,5$. Time integration is performed relying on a third order accurate six stages ESDIRK scheme and using 320 uniform time steps. Note that, since the analytical pressure has a trigonometric spatial behavior, $k+1$ convergence rates for the pressure error in $L^{2}$-norm are attained at all polynomial degrees. Interestingly, the velocity and density errors in $L^{2}$-norm tabulated for $k=2$ are grid independent and allow to appreciate the temporal accuracy of the scheme. Indeed, since the by product between the density and the velocity is bilinear is $\bm{x}$, $k=2$ computations provide machine precision spatial accuracy for both the density and the velocity variables. This observation confirms pressure-robustness of the formulation, note that the velocity error is clearly independent from the pressure error.
The numerical results for $k=1$ show superconvergence of the velocity and density errors, in particular fourth order rate of error reduction is observed until spatial accuracy is on pair with temporal accuracy, resulting in error stagnation. The velocity error is slightly suboptimal at $k=0$. Nevertheless, these latter observations regarding spatial convergence rates must be carefully evaluated taking into account the very peculiar nature of the manufactured analytical solution.
$\mathfrak{E}\,{u_{h}}$ rate $\mathfrak{E}\,{\nabla u_{h}}$ rate $\mathfrak{E}\,{\rho_{h}}$ rate $\mathfrak{E}\,{p_{h}}$ rate $\mathfrak{E}\,{\nabla{\cdot}u_{h}}$
$k=0$,   $\operatorname{card}(\mathcal{T}_{h})=\left\{32,128,512,2048,8192,32768\right\}$
0.00398 0.0141 0.0917 0.0266 5.5e-16
0.00165 1.27 0.00580 1.28 0.0473 0.95 0.0146 0.86 4.7e-16
0.000774 1.09 0.00266 1.12 0.0244 0.96 0.00768 0.93 4.8e-16
0.000379 1.03 0.00129 1.05 0.0124 0.97 0.00393 0.97 4.9e-16
0.000188 1.01 0.000636 1.02 0.00633 0.98 0.00198 0.98 4.8e-16
6.31e-05 1.58 0.000190 1.74 0.00223 1.50 0.00106 0.90 4.9e-16
$k=1$,   $\operatorname{card}(\mathcal{T}_{h})=\left\{32,128,512,2048,8192,32768\right\}$
1.94e-06 7.08e-05 3.89e-07 0.00178 4.6e-16
1.23e-07 3.97 8.95e-06 2.98 1.72e-08 4.50 0.000443 2.00 5.6e-16
7.81e-09 3.99 1.12e-06 3.00 1.11e-09 3.95 0.000111 2.00 5.5e-16
5.19e-10 3.91 1.40e-07 3.00 2.53e-10 2.13 2.77e-05 2.00 5.5e-16
1.88e-10 1.46 1.78e-08 2.98 2.51e-10 0.01 6.93e-06 2.00 5.6e-16
1.86e-10 0.01 7.82e-09 1.19 2.71e-10 -0.11 1.20e-06 2.53 5.7e-16
$k=2$,   $\operatorname{card}(\mathcal{T}_{h})=\left\{32,128,512,2048,8192,32768\right\}$
2.04e-10 2.50e-09 1.79e-10 2.81e-05 6.9e-16
1.89e-10 3.40e-09 2.05e-10 3.57e-06 2.98 7.0e-16
1.87e-10 3.18e-09 2.33e-10 4.48e-07 2.99 7.0e-16
1.86e-10 3.10e-09 2.53e-10 5.63e-08 2.99 7.2e-16
1.86e-10 3.10e-09 2.69e-10 8.11e-09 2.79 7.0e-16
1.86e-10 3.10e-09 3.07e-10 4.42e-09 0.88 6.9e-16
Table 2: Guermond and Quartapelle unsteady test case: spatial convergence rates for $k=0,1,2$ HHO formulations over a $h$-refined triangular elements mesh sequence. Time integration is conducted with a high-order accurate ESDIRK scheme, the time step is selected such that spatial errors dominate.
Temporal convergence results are reported in Table 3 considering two ESDIRK schemes: i) the second order accurate three stages, so-called TR-BDF2, scheme, see [23] for details, ii) the fourth order accurate six stages scheme proposed in [22]. Both ESDIRK schemes are considered in combination with $k=2$ and $k=6$ HHO formulations on a triangular elements mesh with $N=8$ and $\operatorname{card}(\mathcal{T}_{h})=128$, the former provides machine precision spatial accuracy for the velocity and density variables while the latter guarantees negligible spatial errors for all variables, including the pressure. Note that, when considering the $k=2$ spatial discretization with third order accurate ESDIRK, the pressure error stagnates while, thanks to pressure-robustness, the velocity and density errors are almost identical to their $k=6$ counterparts.
Second order accuracy is achieved for all variables when considering the three stages ESDIRK scheme. The six stages ESDIRK scheme provides fourth order accuracy for the density, third order accuracy for the velocity and the pressure and something in between, around three and a half order temporal accuracy, for the velocity gradients. Overall the numerical results are satisfactory and the convergence rates are comparable with the outcomes of Cai et al. [10], a survey where several ESDIRK methods, including those employed in this work, where put to the test. We remark that, in order to achieve higher than second order temporal convergence rates, we rely on Runge-Kutta schemes especially conceived for index two Differential-Algebraic Equation systems (DAEs). This latter choice is mandatory when dealing with non-passive unsteady boundary conditions.
3 stages ESDIRK
k $\frac{7/4\pi}{\delta t}$ $\mathfrak{E}\,{u_{h}}$ rate $\mathfrak{E}\,{\nabla u_{h}}$ rate $\mathfrak{E}\,{\rho_{h}}$ rate $\mathfrak{E}\,{p_{h}}$ rate
6 20 0.000614 0.00161 0.00202 0.00168
40 0.000147 2.06 0.000389 2.06 0.000509 1.99 0.000404 2.07
80 3.62e-05 2.03 9.53e-05 2.03 0.000127 2.00 9.74e-05 2.04
160 8.98e-06 2.01 2.36e-05 2.01 3.19e-05 1.99 2.39e-05 2.02
320 2.23e-06 2.01 5.87e-06 2.01 7.98e-06 2.00 5.93e-06 2.01
2 20 0.000614 0.00161 0.00201 0.00168
40 0.000147 2.06 0.000388 2.06 0.000501 2.01 0.000401 2.07
80 3.62e-05 2.03 9.53e-05 2.03 0.000124 2.01 9.74e-05 2.04
160 8.98e-06 2.01 2.36e-05 2.01 3.10e-05 2.00 2.42e-05 2.01
320 2.23e-06 2.01 5.87e-06 2.01 7.73e-06 2.00 6.92e-06 1.81
6 stages ESDIRK
k $\frac{7/4\pi}{\delta t}$ $\mathfrak{E}\,{u_{h}}$ rate $\mathfrak{E}\,{\nabla u_{h}}$ rate $\mathfrak{E}\,{p_{h}}$ rate $\mathfrak{E}\,{\rho_{h}}$ rate
6 20 3.50e-06 1.26e-05 1.14e-05 1.03e-05
40 2.83e-07 3.63 1.43e-06 3.14 7.87e-07 3.87 1.22e-06 3.08
80 2.34e-08 3.60 1.74e-07 3.04 5.15e-08 3.93 1.66e-07 2.88
160 2.03e-09 3.53 2.26e-08 2.94 3.51e-09 3.88 2.50e-08 2.74
320 1.86e-10 3.45 3.11e-09 2.86 2.57e-10 3.77 4.06e-09 2.62
2 20 3.51e-06 1.26e-05 1.01e-05 1.10e-05
40 2.83e-07 3.63 1.43e-06 3.14 6.93e-07 3.87 3.78e-06 1.54
80 2.34e-08 3.60 1.77e-07 3.02 4.70e-08 3.88 3.58e-06 0.08
160 2.04e-09 3.52 2.40e-08 2.88 3.14e-09 3.90 3.57e-06 0.00
320 1.89e-10 3.43 3.40e-09 2.82 2.05e-10 3.93 3.57e-06 0.00
Table 3: Guermond and Quartapelle unsteady test case: temporal convergence rates for $k=2,6$ HHO formulations over a triangular elements mesh with $\operatorname{card}(\mathcal{T}_{h})=128$. Time integration is conducted with high-order accurate three and six stages ESDIRK schemes with increasingly small time steps, in particular $\delta t=\frac{7/4\pi}{N}$ with $N=20,40,80,160,320$.
4.3 Rayleigh-Taylor instability
In order apply the ESDIRK-HHO formulation to a realistic variable density incompressible flow problem we consider a low Atwood number Rayleigh-Taylor Instability (RTI) configuration. The test case, originally proposed by Tryggvason [24], is setup as proposed by Guermond et al. [16].
We consider the time evolution of two layers of immiscible fluids initially at rest in the rectangular domain $\Omega=(-\frac{d}{2},\frac{d}{2})\times(-2d,2d)$, where $d$ is the reference length. Evolution of the flow field is driven by the downward gravity field as the heavy and the light fluids occupy the top and the bottom layers, respectively. Denoting by $\rho_{\mathrm{t}}$ and $\rho_{\mathrm{b}}$ the density of the top and bottom fluids, respectively, the density ratio $\frac{\rho_{\rm{t}}}{\rho_{\rm{b}}}$ is set to 3, hence the Atwood number, defined as $\mathrm{At}=(\rho_{\mathrm{t}}-\rho_{\mathrm{b}})/(\rho_{\mathrm{t}}+\rho_{\mathrm{b}})$, is equal to one half [24]. Following [16], a smooth transition between the two fluids is guaranteed relying on the formula
$\frac{\rho(\bm{x},t=0)}{\rho_{\mathrm{b}}}=2+\mathrm{tanh}\left(\frac{\bm{x}\cdot\bm{j}-\eta(\bm{x})}{0.01d}\right),$ (22)
where $\eta(\bm{x})=-0.1d\,\mathrm{cos}\left(\frac{2\pi}{d}\bm{x}\cdot\bm{i}\right)$ is the initial position of the perturbed interface between the two fluids. The unit vectors $\bm{i}$ and $\bm{j}$ have horizontal orientation pointing right and vertical orientation pointing up, respectively. The $L^{2}$ projection of the analytical density distribution in Eq. (22) provides the initial density field and the source term in the momentum equation reads $\bm{f}=-\rho g\bm{j}$.
The Reynolds number is defined as $\mathrm{Re}=\frac{\rho_{\mathrm{b}}d^{\frac{3}{2}}g^{\frac{1}{2}}}{\mu}$, note that $d^{\frac{1}{2}}g^{\frac{1}{2}}$ is the reference velocity. Assuming the symmetry of the initial condition is maintained during the time evolution, the computational domain has been restricted to $\Omega=(0,\frac{d}{2})\times(-2d,2d)$ and discretized with a triangular elements mesh with $\operatorname{card}(\mathcal{T}_{h})=65~536$. The computational mesh is obtained splitting in two, along the diagonal, the quadrilateral elements of a $64{\times}512$ cartesian grid. Symmetry boundary conditions, implemented as homogeneous Dirichlet and Neumann BCs in the normal and tangential directions, respectively, are imposed on the computational domain boundary $\partial\Omega$. The polynomial degree of the HHO discretization is $k=6$ and time integration is conducted by means of the second order accurate three stages TR-BDF2 ESDIRK scheme, see [23], within the interval $(0,4]$ discretized using 16001~600 uniform time steps.
In practice we consider unit reference lenght, unit gravity and unit $\rho_{\mathrm{b}}$, so that also the reference velocity is equal to one and $\rho_{\mathrm{t}}=3$. Thus, we set $\mu=\frac{1}{\mathrm{Re}}$ and we run computations considering two Reynolds numbers: $\mathrm{Re}{=}1~000$ and $\mathrm{Re}{=}5~000$. The time evolution of the density field is shown in Figures 1 and 2 at times $t_{\mathrm{Trg}}\simeq 1,1.5,1.75,2,2.25,2.5,2.75$ in the time scale of Tryggvason, where $t_{\mathrm{Trg}}=t\sqrt{\mathrm{At}}$. In practice, for a given a value of $t_{\mathrm{Trg}}$, we select the closest available time point among the 16001~600 available by rounding off to the closest integer the quantity $\frac{t_{\mathrm{Trg}}}{\sqrt{\mathrm{At}}\,\delta t}$, where $\delta t=\frac{4}{1~600}$.
The density field at $\mathrm{Re}{=}1~000$ is in very good agreement with the results of Guermond et al. [17]. As opposite, moving to $\mathrm{Re}{=}5~000$, our computations show more wiggles at the interface between the two fluids, especially at early stages of the flow field evolution. Moreover, the peripheral curling filaments interact with the bulk of the descending plume, a phenomena which does not take place in our results of reference [17]. We remark that the HHO-ESDIRK formulation presented in this work lacks spurious oscillations control while, on the contrary, in [17], an entropy viscosity stabilization was employed. Accordingly, it is somehow expected that the highest Reynolds number computation, associated with a sharper interface featuring strong variations of all the relevant physical quantities, is in worse agreement.
Interestingly, the results at $\mathrm{Re}{=}5~000$ are in better agreement with those those obtained with a high-order Discontinuous Galerkin formulation lacking oscillation control in Bassi et al. [3]. The agreement is quite strong for $t\leq 2$ while the latter stages of time evolution show a different behavior. This confirm the statement that an accurate and detailed prediction of the RTI flow for $t\geq 1.5$ and and $\mathrm{Re}\geq 5000$ is a difficult task [16].
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 1: RTI at $\mathrm{At}=0.5$ and $\mathrm{Re}=1~000$ . From left to right, evolution of the density field at selected time points $t_{\mathrm{Trg}}=t\sqrt{\mathrm{At}}\simeq 1,1.5,1.75,2,2.25,2.5,2.75$.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: RTI at $\mathrm{At}=0.5$ and $\mathrm{Re}=5~000$. From left to right, evolution of the density field at selected time points $t_{\mathrm{Trg}}=t\sqrt{\mathrm{At}}\simeq 1,1.5,1.75,2,2.25,2.5,2.75$.
5 Conclusions
We devised and numerically validated an ESDIRK-HHO formulation of the incompressible Navier–Stokes equations with variable density that is capable of simulating mixtures of immiscible fluids reaching high-orders of accuracy in both space and time. Interestingly, the role of the pressure at the continuous level is closely replicated at the discrete level. Indeed, in the one hand pressure is the Lagrange multiplier of the incompressibilty constraint and, on the other hand, the hybrid pressure space enforces $\bm{H}(\operatorname{div})$-conformity starting from non-conforming polynomial spaces. The distinguishing features, that is pressure-robustness, exact incompressibility and pure advection of the density, have been numerically demonstrated.
Since the spatial discretization requires primitive variables in order to achieve $\bm{H}(\operatorname{div})$-conformity, while the time derivatives are formulated in terms of conservative variables, the ESDIRK time marching strategy has been carefully designed. The resulting scheme is fully implicit and fully hybrid, and static condensation of element unknowns has been employed to alleviate the memory footprint of the Jacobian matrix. To further improve the performance of iterative linear solvers a $p$-multigrid preconditioner has been fruitfully exploited when $k>0$. Note that, thanks to the use of non-conforming polynomial spaces, inherited coarse grid operators are trivial to construct. The ability of simulating realistic flow configurations is demonstrated tackling the Rayleigh-Taylor instability. Notably, high-order computations are carried out without oscillation control and the evolution of the density field is in good agreement with previously published works.
Further research efforts are required to target physically relevant multi-phase flows and higher Atwood numbers flow configurations. In particular, future works will consider the extension of the HHO-ESDIRK formulation to the Cahn-Hilliard-Navier-Stokes (CHNS) system.